Consider the perfectly competitive market for Diesel. The aggregate demand for gasoline is
The aggregate demand for gasoline is
Q_d=100-p
While the aggregate supply is Q_s= 3p
Work out the equilibrium price and quantity.
At the equilibrium level established in part a), calculate the consumer surplus, producer surplus and total surplus.
1
Expert's answer
2021-09-06T07:06:58-0400

Given

"Qd=100-p\\Qs=3P"

At equilibrium Qs=Qd

"100-p=3p\\100=3p+p\\100=4p\\p=25\\Q=3p\\Q=3(25)\\Q=75"

Consumer surplus

"Qd=100-p\\0=100-p\\p=100\\\frac{1}{2}×75×(100-25)\\\frac{1}{2}×75×75\\=2812.5"


Producer surplus

"Qs=3p\\0=3p\\p=0\\\frac{1}{2}×75×(25-0)\\\frac{1}{2}×75×25\\=900"

Total surplus=producer surplus + consumer surplus

"=900+2812.5\\=3712.5"